OpenStudy (anonymous):

86 is 45% of what number? (1 point)

OpenStudy (anonymous):

86=.45x, solve for x

OpenStudy (anonymous):

\[86 \times \frac{45}{100} = ?\] the question mark will be ur answer.

OpenStudy (anonymous):

JayDS you are wrong

OpenStudy (anonymous):

@math_proof you are wrong, have u even tried ur method?

OpenStudy (anonymous):

Ok who's is wrong and who's is right???

OpenStudy (anonymous):

using ur method u will get the answer to be 191.11111 which is impossible and 100% wrong.

OpenStudy (anonymous):

trust me @ktlyn9 I am right.

OpenStudy (anonymous):

I can guarantee it, you can even ask others to come and check if you do not believe.

OpenStudy (anonymous):

Ok..I see what u mean. Thanks.

OpenStudy (anonymous):

191.11111 then when you multiply it by .45 = 86:p

OpenStudy (anonymous):

JayDS why would you multiply 86 x .45? lol it doesn't make any sense

OpenStudy (anonymous):

86 is bla bla bla means 86= to something (is always means equal) and even reading further it says 45% of what number which means .45 * some number

OpenStudy (anonymous):

ok, I cannot explain your "Wrong method" to you but lets think logically 45% of something? we want 45% of 86 and 50% would be just half of it so 43, do u get this? so say u are correct, we use ur method then and we get the number to be 191.111 how the hell is that possible?

OpenStudy (anonymous):

191.111 = more than 200% of 86.

OpenStudy (anonymous):

i know the answer is stupid but thats the correct answer :)

OpenStudy (anonymous):

JayDS uuu are sooo wrong:p

OpenStudy (anonymous):

u can google that :p to find an answer:p

OpenStudy (anonymous):

@math_proof u have been wrong many times before and you clearly have no logic behind ur answer, anyone would laugh if u told them ur answer to this question.

OpenStudy (anonymous):

whatever @math_proof said isnt even a choice....

OpenStudy (anonymous):

what are the choices?

OpenStudy (anonymous):

JayDS google it, and don't says something you don't know about me

OpenStudy (anonymous):

@ktlyn9 what are the choices?

OpenStudy (anonymous):

86 = 0.45n n = 0.45 × 86 0.45 = n × 86 86 = 45n @math_proof

OpenStudy (anonymous):

@ktlyn9 why did you say there is no answer like mine? as there is one like mine

OpenStudy (anonymous):

What one is it? you said like 1119 or something!

OpenStudy (anonymous):

Oh....I see it...oops. my fault xD